OpenStudy (anonymous):

Simplify: (–5) • (7) – (6) • (–2)

OpenStudy (anonymous):

-35-12=-47

OpenStudy (anonymous):

I thought it was bodmass, -35 - -12 = 23

OpenStudy (anonymous):

ok now im confused

OpenStudy (anonymous):

sorry, i may be wrong, often am but, i thought you should always times and divide before you add or subtract so:\[-5 \times 7 = -35\] and \[6 \times -2 = -12\] leaving\[(-35) - (-12) = 23\]